OpenStudy (anonymous):

A certain store is selling an $80 radio for $64. If a different radio had a list price of $200 and was discounted at 1 1/2 times the percent discount on the $80 model, what would its selling price be? 1. $90 2. $105 3. $120 4. $140 5. $160

OpenStudy (a_clan):

First radio, discount% = [(80-64)/80 ] *100 = 20% SP of second radio = 200 - (3/2)(20/100)(200) = $140